OpenStudy (sambot):

Which of the given pairs of expressions are equivalent? A. 3z+2z and 5z B. (8x-8x)-y and 0 C. 9.1+4.5a-2.5a and 11.1a D. -7y(3)-3 and 0 E. y-6(3) and y-18 HELP

OpenStudy (serenity74):

Which answers can you cancel out right off of the bat?

RhondaSommer (rhondasommer):

well solve each expression :)

OpenStudy (sambot):

um...

OpenStudy (sambot):

ok

RhondaSommer (rhondasommer):

what is 3z+2z?

OpenStudy (sambot):

5z?

RhondaSommer (rhondasommer):

yup so that one is equivalent

OpenStudy (sambot):

k

RhondaSommer (rhondasommer):

sorry (8x-8x)

OpenStudy (sambot):

0x

OpenStudy (sambot):

?

RhondaSommer (rhondasommer):

yup so when you multiply anything by 0 it equals 0 so its just 0-y

OpenStudy (sambot):

so it is not equivalent?

RhondaSommer (rhondasommer):

nope :)

OpenStudy (sambot):

k

RhondaSommer (rhondasommer):

what is 4.5a-2.5a and

OpenStudy (sambot):

give me a second please

OpenStudy (sambot):

2.0a?

RhondaSommer (rhondasommer):

also- is it 9.1a or just 9.1 and yes that is correct

OpenStudy (sambot):

equivalent?

RhondaSommer (rhondasommer):

we dont know that yet. is it 9.1 or 9.1a

OpenStudy (sambot):

9.1

OpenStudy (manvirk21):

It would be E. because 6*3=18 In conclusion y-18=y-18

RhondaSommer (rhondasommer):

lol goodjob you have to pick all of them. AND you shouldnt give immediate answers anyway

RhondaSommer (rhondasommer):

then should the equivalent have 9.1 in it?

OpenStudy (sambot):

i dont know

OpenStudy (sambot):

um i think so

RhondaSommer (rhondasommer):

it should because it has no (a) beside it

OpenStudy (sambot):

ok

RhondaSommer (rhondasommer):

-7y(3)-3 so fir this you know that you need to multiply then subtract the 3 so what ever the answer is it should at least have a (y) in it. :)

OpenStudy (sambot):

k thatnks

OpenStudy (sambot):

thanks*

RhondaSommer (rhondasommer):

y-6(3) then you just simplify that

OpenStudy (sambot):

i have to go..thanks for the help